A chemist needs to create a solution consisting of 70% hydrochloric acid (the other 30% is water). How much pure acid (100% acid solution) should be mixed with 11 mL of a solution with 40% acid in order to get a solution with 70% acid?
Answer by TimothyLamb(4379)

x = ml of 100% acid
---
0.04*11 + 1.00x = 0.70*(x + 11)
0.04*11 + 1.00x = 0.70x + 0.70*11
x - 0.70x = 0.70*11 - 0.04*11
0.30x = 0.70*11 - 0.04*11
x = (0.70*11 - 0.04*11)/0.30
x = 24.2
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---